# BA-LoRA: Bias-Alleviating Low-Rank Adaptation to Mitigate Catastrophic Inheritance in Large Language Models

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable proficiency across various natural language processing (NLP) tasks. However, adapting LLMs to downstream applications requires computationally intensive and memory-demanding fine-tuning procedures. To alleviate these burdens, parameter-efficient fine-tuning (PEFT) techniques have emerged as a promising approach to tailor LLMs with minimal computational overhead. While PEFT methods offer substantial advantages, they do not fully address the pervasive issue of bias propagation from pre-training data. This work introduces Bias-Alleviating Low-Rank Adaptation (BA-LoRA), a novel PEFT method designed to counteract bias inheritance. BA-LoRA incorporates three distinct regularization terms: (1) a consistency regularizer, (2) a diversity regularizer, and (3) a singular value decomposition regularizer. These regularizers aim to enhance the models’ consistency, diversity, and generalization capabilities during fine-tuning. We conduct extensive experiments on natural language understanding (NLU) and natural language generation (NLG) tasks using prominent LLMs such as LLaMA, Mistral, and Gemma. The results demonstrate that BA-LoRA outperforms LoRA and its state-of-the-art variants. Moreover, our method effectively mitigates the adverse effects of pre-training bias, leading to more reliable and robust model outputs. The code is available at https://github.com/cyp-jlu-ai/BA-LoRA.
@@ 1 Introduction
To mitigate the detrimental effects of Catastrophic Inheritance, particularly noise and imbalance, we propose Bias-Alleviating Low-Rank Adaptation (BA-LoRA). Building upon Principal Singular Values and Singular Vectors Adaptation (PiSSA) (Meng et al., 2024), which addresses convergence issues in standard LoRA, our approach incorporates three distinct regularization terms: a consistency regularizer, a diversity regularizer, and a singular value decomposition (SVD) regularizer. The consistency regularizer preserves valuable pre-trained knowledge during fine-tuning, while the diversity regularizer encourages varied model outputs. The SVD regularizer enhances the generalization capabilities of generative models. Recognizing the fundamental differences between Natural Language Understanding (NLU) and Natural Language Generation (NLG), such as determinism in NLU versus diversity in NLG, we tailor our regularization strategies accordingly.
To evaluate the efficacy of BA-LoRA, we conduct comprehensive experiments across diverse benchmarks, including mathematical reasoning (GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023)), coding (HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)), natural language understanding (GLUE (Wang et al., 2018)), and general language evaluation (MT-Bench (Zheng et al., 2024)). Our experiments utilize prominent LLMs such as LLaMA 2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Gemma-7B (Team et al., 2024), as well as encoder-only architectures like RoBERTa-large (Liu et al., 2019) and DeBERTa-v3-base (He et al., 2021b). The results unequivocally demonstrate BA-LoRA’s superiority over LoRA and PiSSA. Moreover, our method effectively attenuates noise inherited from pre-training, leading to more robust and generalizable models.
@@ 2 Related Works
Parameter-efficient fine-tuning (PEFT) techniques (Xu et al., 2023b; Han et al., 2024) have garnered significant attention as an approach to adapting LLMs for specific tasks under limited hardware resources. Three main categories of PEFT techniques are commonly used. The first category includes adapter-based methods (Houlsby et al., 2019b; Lin et al., 2020; Lei et al., 2023; He et al., 2021a), which introduce additional layers into the model and fine-tune these layers (typically with far fewer parameters) to reduce computational costs. The second category comprises soft prompt tuning methods (Hambardzumyan et al., 2021; Lester et al., 2021; Li & Liang, 2021a; Liu et al., 2023), which prepend learnable soft prompts to the model’s input to tailor it to specific tasks. These methods leverage the inherent capabilities of pre-trained models, requiring only the appropriate prompts to adapt to downstream tasks. The third category encompasses low-rank adaptation (LoRA) and its variants (Hu et al., 2021; Zhang et al., 2022; Dettmers et al., 2024). LoRA introduces the product of low-rank matrices within existing layers to approximate weight updates during fine-tuning (Hu et al., 2021).
Variants of LoRA enhance its efficiency and performance in different ways. AdaLoRA adaptively distributes the parameter budget among weight matrices based on their importance, improving efficiency and performance by pruning unimportant updates and minimizing computational overhead (Zhang et al., 2022). DoRA increases LoRA’s learning capacity and stability by decomposing pre-trained weights into magnitude and direction components for fine-tuning (Liu et al., 2024). LoHA enhances LoRA by employing Hamiltonian products (Hyeon-Woo et al., 2021). DyLoRA addresses the fixed size and rank optimization limitations of LoRA by dynamically training LoRA blocks across varying ranks (Valipour et al., 2022). DeltaLoRA improves the representational capacity of LoRA by updating the model’s original weights using parameters from adapter layers (Zi et al., 2023). PiSSA initializes adapter matrices $A$ and $B$ to approximate the original matrix $W$ through singular value decomposition, leading to faster convergence and improved performance (Meng et al., 2024). While many LoRA variants focus on accelerating convergence or reducing memory consumption, our BA-LoRA method uniquely addresses the core challenge of Catastrophic Inheritance in LLM fine-tuning.
@@ 3.2.3 Regularizations for NLG Tasks
Singular Value Decomposition Regularization To enhance the generalization capability of generative models, we introduce a regularization technique that accentuates the most significant singular values, enabling the model to capture the principal components of the data and prioritize the most informative patterns.
$\mathcal{L}_{\mathrm{SVDR\_NLG}}=-\frac{\sum_{i=1}^{k}\sigma_{i}}{\sum_{j=1}^{D}\sigma_{j}}$ (12)
where $\sigma_{i}$ denotes the $i$-th largest singular value, $\sigma_{j}$ represents the $j$-th singular value, and $D$ is the total number of singular values. This regularization term aims to maximize the relative contribution of the top $k$ singular values, encouraging the model to focus on the most critical aspects of the data. By integrating this regularization, the model is steered towards generating higher-quality and more diverse outputs, ultimately improving its performance and robustness.
@@ 4 Experiments
This section presents a comprehensive evaluation of our proposed BA-LoRA method across a diverse range of natural language generation (NLG) and natural language understanding (NLU) benchmarks. Our results unequivocally demonstrate the superiority of BA-LoRA over existing LoRA variants. Furthermore, through rigorous experimentation, we elucidate BA-LoRA’s efficacy in mitigating the adverse impacts of noisy data, thereby enhancing model robustness and generalizability.
@@ 4.2 Implementation Details
In our experiments, we adopt the PiSSA (Meng et al., 2024) implementation strategy. We compute the loss using only the responses from the instruction-following dataset, ensuring lora_dropout to 0. We utilize the Float32 computation type for both the base model and the adapter in BA-LoRA. For the NLU tasks, we set the hyperparameters as: $\lambda_{1}=1e-4$, $\lambda_{2}=4e-4$, and $\lambda_{3}=1e-4$. We set lora_r $=$ lora_alpha $=128$ and use AdamW (Loshchilov & Hutter, 2017) optimizer with a batch size of 128, a learning rate of $52e-$ 5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. For the NLG tasks, the hyperparameters are set as: $\lambda_{1}=1e-4$, $\lambda_{2}=3e-4$, and $\lambda_{3}=1e-4$. We set lora_r as 8 and select lora_alpha in 8, 16. We utilize AdamW with a linear learning rate schedule to optimize and tune the learning rate (LR) from $41e-$ 4, $42e-$ 4, $43e-$ 4, $44e-$ 4, $45e-$ 4, $46e-$ 4, $55e-$ 5, $53e-$ 5. Batch sizes (BS) are selected from 6, 8, 16, 32. Appendix Section B presents the detailed hyperparameters we utilized on the GLUE benchmark. comparison. All experiments were conducted using NVIDIA A40 (48G) GPUs. All presented results are derived from three independent experiments, ensuring the consistency and robustness of our findings. Each experiment was conducted under identical conditions, and the results were averaged to mitigate any variability. We enhanced $\mathcal{L}_{\mathrm{SVDR\_NLG}}$’s computational efficiency by using partial SVD to compute only the top $k$ singular values, reducing overhead.
@@ 4.3 Results and Analysis
LLaMA-2-7B Full FT 6738M 49.05 7.22 21.34 35.59 4.91 23.62
LoRA 320M 42.47 5.60 17.03 31.48 4.62 20.24
PiSSA 320M 52.01 7.76 21.55 33.09 4.87 23.86
BA-LoRA 320M 53.83 9.13 23.58 36.86 5.11 25.70
Mistral-7B Full FT 6738M 67.02 18.60 45.12 51.38 4.95 37.41
LoRA 168M 67.68 19.90 42.54 55.74 4.92 38.16
PiSSA 168M 71.90 21.72 45.20 60.83 5.23 40.98
BA-LoRA 168M 73.04 22.11 46.31 62.29 5.41 41.83
Gemma-7B Full FT 6738M 71.34 22.74 46.95 55.64 5.40 40.41
LoRA 200M 74.64 31.16 51.64 62.84 5.01 45.06
PiSSA 200M 77.58 31.47 53.15 65.49 5.66 46.67
BA-LoRA 200M 78.13 32.25 54.41 66.12 5.73 47.33
@@
Full FT 184M 89.90 95.61 89.50 69.23 94.09 92.44 83.85 91.71 88.29
BitFit 0.1M 89.37 94.84 87.75 66.96 92.24 88.41 78.70 91.35 86.20
HAdapter 1.22M 90.13 95.53 89.95 68.64 94.11 91.91 84.48 91.48 88.28
PAdapter 1.18M 90.33 95.61 89.46 68.77 94.29 92.04 85.20 91.60 88.41
LoRA 1.33M 90.71 94.79 89.85 70.05 93.94 92.07 85.43 91.67 88.56
LoHA 1.33M 90.74 94.92 90.43 70.63 93.95 92.05 86.41 91.72 88.86
DoRA 1.27M 90.29 95.79 90.93 70.85 94.10 92.17 86.04 91.79 89.00
AdaLoRA 1.27M 90.76 96.10 90.69 71.45 94.55 92.23 87.59 91.84 89.40
PiSSA 1.33M 90.47 95.81 91.48 72.27 94.41 92.21 87.14 91.93 89.47
BA-LoRA 1.33M 90.92 96.25 91.83 72.79 94.84 92.59 87.87 92.15 89.91
@@ 4.3.1 Analysis of the NLG and NLU Performance of BA-LoRA
To evaluate BA-LoRA’s effectiveness on NLG tasks, we fine-tuned LLaMA-2-7B, Mistral-7B, and Gemma-7B on the MetaMathQA dataset (Yu et al., 2023) and assessed their mathematical problem-solving capabilities using the GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023) validation sets, reporting Accuracy. Similarly, models were fine-tuned on the CodeFeedback dataset (Zheng et al., 2024) and evaluated for coding proficiency via HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), with PASS@1 metrics reported. To assess conversational abilities, models were trained on the WizardLM-Evol-Instruct dataset (Xu et al., 2024) and evaluated on MT-Bench (Zheng et al., 2024), with response quality judged by GPT-4 and first turn scores reported. All experiments utilized 100K data points and a single training epoch for efficiency.
@@
To assess the effectiveness of BA-LoRA on natural language understanding (NLU) tasks, we conducted experiments on the GLUE benchmark (Wang et al., 2018), which includes two single-sentence classification tasks (CoLA, SST), five paired-text classification tasks (MNLI, RTE, QQP, MRPC, QNLI), and one text similarity prediction task (STS-B). The evaluation metrics comprise the overall matched and mismatched accuracy for MNLI, the Matthews correlation coefficient for CoLA, the Pearson correlation coefficient for STS-B, and accuracy for the remaining tasks. We used the DeBERTa-v3-base model (He et al., 2021b) and compared BA-LoRA against ten baseline methods, including Full Fine-Tuning (Full FT), BitFit (Zaken et al., 2021), HAdapter (Houlsby et al., 2019a), PAdapter (Pfeiffer et al., 2020), LoRA (Hu et al., 2021), LoHA (Hyeon-Woo et al., 2021), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2022), and PiSSA (Meng et al., 2024).

### v4

@@ Abstract
Large language models (LLMs) have demonstrated remarkable proficiency across various natural language processing (NLP) tasks. However, adapting LLMs to downstream applications requires computationally intensive and memory-demanding fine-tuning procedures. To alleviate these burdens, parameter-efficient fine-tuning (PEFT) techniques have emerged as a promising approach to tailor LLMs with minimal computational overhead. While PEFT methods offer substantial advantages, they do not fully address the pervasive issue of bias propagation from pre-training data. This work introduces Bias-Alleviating Low-Rank Adaptation (BA-LoRA), a novel PEFT method designed to counteract bias inheritance. BA-LoRA incorporates three distinct regularization terms: (1) a consistency regularizer, (2) a diversity regularizer, and (3) a singular value decomposition regularizer. These regularizers aim to enhance the models’ consistency, diversity, and generalization capabilities during fine-tuning. We conduct extensive experiments on natural language understanding (NLU) and natural language generation (NLG) tasks using prominent LLMs such as LLaMA, Mistral, and Gemma. The results demonstrate that BA-LoRA outperforms LoRA and its state-of-the-art variants. Moreover, the extended experiments demonstrate that our method effectively mitigates the adverse effects of pre-training bias, leading to more reliable and robust model outputs. The code is available at https://github.com/cyp-jlu-ai/BA-LoRA.
@@ 1 Introduction
To mitigate the detrimental effects of Catastrophic Inheritance, particularly noise, and imbalance, we propose Bias-Alleviating Low-Rank Adaptation (BA-LoRA). Building upon Principal Singular Values and Singular Vectors Adaptation (PiSSA) (Meng et al., 2024), which addresses convergence issues in standard LoRA, our approach incorporates three distinct regularization terms: a consistency regularizer, a diversity regularizer, and a singular value decomposition (SVD) regularizer. The consistency regularizer preserves valuable pre-trained knowledge during fine-tuning, while the diversity regularizer encourages varied model outputs. The SVD regularizer enhances the generalization capabilities of generative models. Recognizing the fundamental differences between Natural Language Understanding (NLU) and Natural Language Generation (NLG), such as determinism in NLU versus diversity in NLG, we tailor our regularization strategies accordingly.
To evaluate the efficacy of BA-LoRA, we conduct comprehensive experiments across diverse benchmarks, including mathematical reasoning (GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023)), coding (HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)), natural language understanding (GLUE (Wang et al., 2018)), and general language evaluation (MT-Bench (Zheng et al., 2024)). Our experiments utilize prominent LLMs such as LLaMA 2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Gemma-7B (Team et al., 2024), as well as encoder-only architectures like RoBERTa-large (Liu et al., 2019) and DeBERTa-v3-base (He et al., 2021b). The results unequivocally demonstrate BA-LoRA’s superiority over LoRA and its state-of-the-art variants. Moreover, the extended experiments demonstrate that our method effectively attenuates noise inherited from pre-training, leading to more robust and generalizable models.
@@ 2 Related Works
Parameter-efficient fine-tuning (PEFT) techniques (Xu et al., 2023b; Han et al., 2024) have garnered significant attention as an approach to adapting LLMs for specific tasks under limited hardware resources. Three main categories of PEFT techniques are commonly used. The first category includes adapter-based methods (Houlsby et al., 2019b; Lin et al., 2020; Lei et al., 2023; He et al., 2021a), which introduce additional layers into the model and fine-tune these layers (typically with far fewer parameters) to reduce computational costs. The second category comprises soft prompt tuning methods (Hambardzumyan et al., 2021; Lester et al., 2021; Li & Liang, 2021a; Liu et al., 2023), which prepend learnable soft prompts to the model’s input to tailor it to specific tasks. These methods leverage the inherent capabilities of pre-trained models, requiring only the appropriate prompts to adapt to downstream tasks. The third category encompasses low-rank adaptation (LoRA) and its variants (Hu et al., 2021; Zhang et al., 2023a; Dettmers et al., 2024). LoRA introduces the product of low-rank matrices within existing layers to approximate weight updates during fine-tuning (Hu et al., 2021).
Variants of LoRA enhance its efficiency and performance in different ways. AdaLoRA adaptively distributes the parameter budget among weight matrices based on their importance, improving efficiency and performance by pruning unimportant updates and minimizing computational overhead (Zhang et al., 2023b). DoRA increases LoRA’s learning capacity and stability by decomposing pre-trained weights into magnitude and direction components for fine-tuning (Liu et al., 2024). LoHA enhances LoRA by employing Hamiltonian products (Hyeon-Woo et al., 2021). DyLoRA addresses the fixed size and rank optimization limitations of LoRA by dynamically training LoRA blocks across varying ranks (Valipour et al., 2022). DeltaLoRA improves the representational capacity of LoRA by updating the model’s original weights using parameters from adapter layers (Zi et al., 2023). PiSSA initializes adapter matrices $A$ and $B$ to approximate the original matrix $W$ through singular value decomposition, leading to faster convergence and improved performance (Meng et al., 2024). While many LoRA variants focus on accelerating convergence or reducing memory consumption, our BA-LoRA method uniquely addresses the core challenge of Catastrophic Inheritance in LLM fine-tuning.
@@ 3.2.3 Regularizations for NLG Tasks
Singular Value Decomposition Regularization To enhance the generalization of generative models while ensuring computational efficiency, we introduce a regularization technique based on randomized SVD (Pasand & Bashivan, ). This method efficiently approximates the dominant singular values of a matrix, significantly reducing the computational and memory costs associated with traditional SVD while preserving key information about the data.
$\mathcal{L}_{\mathrm{SVDR\_NLG}}=-\frac{\sum_{i=1}^{k}\tilde{\sigma}_{i}}{\sum_{j=1}^{D}\tilde{\sigma}_{j}}$ (12)
Here, $\tilde{\sigma}_{i}$ represents the $i$-th largest approximated singular value obtained using randomized SVD, and $D$ is the total number of singular values in the approximation. This regularization maximizes the relative contribution of the top $k$ singular values, encouraging the model to prioritize the most informative patterns in the data.
@@ 4 Experiments
This section presents a comprehensive evaluation of our proposed BA-LoRA method across a diverse range of NLG and NLU benchmarks. Our results unequivocally demonstrate the superiority of BA-LoRA over existing LoRA variants. Furthermore, through rigorous experimentation, we elucidate BA-LoRA’s efficacy in mitigating the adverse impacts of noisy data, thereby enhancing model robustness and generalizability.
@@ 4.2 Implementation Details
In our experiments, we adopt the PiSSA (Meng et al., 2024) implementation strategy. We compute the loss using only the responses from the instruction-following dataset, ensuring lora_dropout to 0. We utilize the Float32 computation type for both the base model and the adapter in BA-LoRA. For the NLU tasks, we set the hyperparameters as: $k=5$, $\lambda_{1}=1e-1$, $\lambda_{2}=4e-2$, and $\lambda_{3}=1e-2$. We set lora_r $=$ lora_alpha $=128$ and use AdamW (Loshchilov & Hutter, 2017) optimizer with a batch size of 128, a learning rate of $52e-$ 5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. For the NLG tasks, the hyperparameters are set as: $k=10$, $\lambda_{1}=1e-4$, $\lambda_{2}=3e-1$, and $\lambda_{3}=1e-1$. We set lora_r as 8 and select lora_alpha in 8, 16. We utilize AdamW with a linear learning rate schedule to optimize and tune the learning rate (LR) from $41e-$ 4, $42e-$ 4, $43e-$ 4, $44e-$ 4, $45e-$ 4, $46e-$ 4, $55e-$ 5, $53e-$ 5. Batch sizes (BS) are selected from 6, 8, 16, 32. Appendix Section B presents the detailed hyperparameters we utilized on the GLUE benchmark. comparison. All experiments were conducted using NVIDIA A40 (48G) GPUs. The results were derived from three independent experiments conducted under identical conditions, which were averaged to ensure consistency, robustness, and variability mitigation.
@@ 4.3 Results and Analysis
LLaMA-2-7B Full FT 6738M 49.13±0.21 7.29±0.22 21.20±0.30 35.59±0.25 4.91±0.01 23.62
LoRA 320M 42.47±0.29 5.60±0.35 17.03±0.61 32.77±0.46 4.62±0.11 20.50
PiSSA 320M 52.37±0.52 7.76±0.19 21.55±0.44 33.09±0.57 4.87±0.06 23.93
BA-LoRA 320M 54.53±0.41 9.21±17 23.58±0.25 36.86±0.31 5.11±0.05 25.86
Mistral-7B Full FT 6738M 69.91±0.25 18.64±0.35 45.31±0.14 51.46±0.13 4.95±0.05 38.05
LoRA 168M 67.68±0.55 19.90±0.25 42.54±0.44 56.85±0.23 4.92±0.07 38.38
PiSSA 168M 72.25±0.64 21.95±0.37 45.37±0.25 61.57±0.44 5.23±0.05 41.27
BA-LoRA 168M 73.17±0.34 22.79±0.56 46.31±0.17 62.77±0.33 5.41±0.06 42.09
Gemma-7B Full FT 6738M 72.09±0.32 22.71±0.34 47.02±0.27 55.67±0.05 5.40±0.12 40.58
LoRA 200M 74.64±0.58 31.16±0.33 51.64±0.28 63.52±0.65 5.01±0.03 45.19
PiSSA 200M 77.58±0.41 31.47±44 53.22±35 65.49±0.18 5.66±0.05 46.68
BA-LoRA 200M 78.13±0.25 32.25±25 54.44±0.15 66.25±0.33 5.73±0.07 47.36
Figure 1: Performance Comparison of Full Fine-Tuning, LoRA, PiSSA, and BA-LoRA Across Different Ranks.
@@
Full FT 184M 90.34±0.18 96.33±0.11 89.95±1.07 71.43±0.72 94.24±0.10 92.11±0.28 83.75±1.81 91.04±0.48 88.86
BitFit 0.1M 89.54±0.29 94.68±0.11 87.95±1.33 67.31±0.49 92.45±0.17 88.72±0.45 79.12±0.39 91.63±0.37 86.43
HAdapter 1.22M 90.23±0.07 95.38±0.06 89.97±0.27 68.73±0.27 94.31±0.29 91.99±0.28 84.76±0.39 91.58±0.13 88.37
PAdapter 1.18M 90.42±0.36 95.49±0.10 89.71±0.35 69.04±0.10 94.38±0.26 92.15±0.43 85.53±0.18 91.69±0.13 88.55
LoRA 1.33M 90.71±0.16 94.79±0.16 89.85±0.21 70.05±0.34 93.94±0.09 92.07±0.48 85.43±0.09 91.67±0.29 88.56
LoHA 1.33M 90.74±0.32 94.92±0.47 90.43±0.48 70.63±0.10 93.95±0.28 92.05±0.09 86.41±0.10 91.72±0.28 88.86
DoRA 1.27M 90.48±0.10 95.85±0.08 91.04±0.15 71.03±0.18 94.21±0.37 92.34±0.16 86.19±0.25 91.92±0.38 89.13
AdaLoRA 1.27M 90.87±0.08 96.18±0.43 90.81±0.40 71.64±0.12 94.68±0.46 92.37±0.35 87.78±0.36 91.97±0.43 89.53
PiSSA 1.33M 90.47±0.44 95.81±0.45 91.48±0.49 72.27±0.29 94.41±0.41 92.21±0.26 87.14±0.08 91.93±0.25 89.47
BA-LoRA 1.33M 90.92±0.38 96.25±0.09 91.83±0.25 72.79±0.42 94.84±0.26 92.59±0.18 87.87±0.31 92.15±0.08 89.91
@@ 4.3.1 Analysis of the NLG and NLU Performance of BA-LoRA
To evaluate BA-LoRA’s effectiveness on NLG tasks, we fine-tuned LLaMA-2-7B, Mistral-7B, and Gemma-7B on the MetaMathQA (Yu et al., 2023) and assessed their mathematical problem-solving capabilities using the GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023) validation sets, reporting Accuracy. Similarly, models were fine-tuned on the CodeFeedback (Zheng et al., 2024) and evaluated for coding proficiency via HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), with PASS@1 metrics reported. To assess conversational abilities, models were trained on the WizardLM-Evol-Instruct (Xu et al., 2024) and evaluated on MT-Bench (Zheng et al., 2024), with response quality judged by GPT-4 and first turn scores reported. All experiments utilized 100K data points and a single training epoch for efficiency.
@@
To assess the effectiveness of BA-LoRA on natural language understanding (NLU) tasks, we conducted experiments on the GLUE benchmark (Wang et al., 2018), which includes two single-sentence classification tasks (CoLA, SST), five paired-text classification tasks (MNLI, RTE, QQP, MRPC, QNLI), and one text similarity prediction task (STS-B). The evaluation metrics comprise the overall matched and mismatched accuracy for MNLI, the Matthews correlation coefficient for CoLA, the Pearson correlation coefficient for STS-B, and accuracy for the remaining tasks. We used the DeBERTa-v3-base model (He et al., 2021b) and compared BA-LoRA against ten baseline methods, including Full Fine-Tuning (Full FT), BitFit (Zaken et al., 2021), HAdapter (Houlsby et al., 2019a), PAdapter (Pfeiffer et al., 2020), LoRA (Hu et al., 2021), LoHA (Hyeon-Woo et al., 2021), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023a), and PiSSA (Meng et al., 2024).
@@
4.4 Evaluating the Performance of Different Ranks
We compare the performance of BA-LoRA, LoRA, and PiSSA at different ranks using LLaMA-2-7B and Mistral-7B-v0.1 models. Each method is fine-tuned for one epoch on the MetaMathQA-100K dataset with ranks ranging from 1 to 128 and evaluated on the GSM8K and MATH. As Figure  1 shows, BA-LoRA consistently outperforms LoRA and PiSSA across all rank settings and datasets. As the rank increases, the performance of BA-LoRA and PiSSA surpasses full parameter fine-tuning. However, BA-LoRA performs better, especially on Mistral-7B-v0.1.